question 19

following a strategy of product differentiation , LUKE company makes a high end appliance  ,AP15, LUKE  company presents the following data for the years 1 and 2

year 1 year 2

units of Ap15  produced and sold 20000 21000

selling price $200 220

direct material (square meters) 60000 61500

direct material costs per liter $20 $22

manufacturing capacity for AP15 (units) 25000 25000

total manufacturing conversion costs $1000000 1100000

manufacturing conversion costs per unit of capacity $40 $44

selling and customer service capacity (customers) 60   58

total selling and customers service costs $360000 362500

cost per customer of selling & customer service capacity 6000 $6250

Luke company produces no defective units but it wants to reduce direct materials usage per unit of AP15 in year 2 . manufacturing conversion costs in each year depend on production capacity defined in terms of AP15 units that can be produced selling and customer service costs depend on the number of customers that the customers and service functions are designed to support .neither conversion costs or customer service costs are affected by changes in actual volume . LUKE company has 46 customers in year 1 and 50 customers in year 2 the industry market size for high end appliances increased 5% from year 1 to year 2

1- what is the LUKE  company revenue effect of growth component?

1-$440000 favaurable

2-$220000 favourable

3-$400000 favourable

4- $400000 unfavourable

5-200000 favourable

2-what is the LUKE company net increase in operating income as a result of the price recovery component

1-$20000 unfavourable

2-$20000 favourable

3-182000 unfavourable

4-179000 favourable

5-179000 unfavourable
Answer #1

1. What is the revenue effect of the growth component?

Answer: 5

Explanation: (21,000 - 20,000) × $200 = $200,000 F

2. What is the net effect on operating income as a result of the price-recovery component?

Answer :4 179000 favourable

Explanation: ($220 - $200) × 21,000 = $420,000 F

[($22 - $20) × 63,000] + [($44 - $40) × 25,000] + [($6,250 - $6,000) × 60] = $241,000 U

$420,000 F + $241,000 U = $179,000 F